what is the minimum value of 2x3-3x2-36x+10. kindly get me the answer atthe earliest
 
 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
Find dy/dx and equate to zero
Hence
6x² -6x -36 = 0
x² - x -6 = 0
Hence
x = 3, -2
Now
Find d²y/dx²
12 x -6
Now at x = 3, it is positive hence it will be minimum at x = 3
Minimum value = 2 *27 - 3 *9 - 36*3 +10 = 37- 108 = -71
 
Regards
Arun (askIITians forum expert)